• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Gao

    Log entry time 00:06:56 on June12,2009

    Entry number 280707

    keyword=Swing shift summary

    5:35 pm MCC callled for beam permit
    6:00 started optics run with raster off runs 3483, 22351, Happex 31516
    6:18 started optics run with raster on runs 3484, 22352
    18:55 Reference hydrogen run with raster on (optics), runs 3486, 22354, New Compton run 17339
    19:35 Controlled access to remove HRS sieves and install BigBite Sieve. Change Right HRS momentum to 2.025 GeV/c
    20:10 Beam back, continue to take H2 reference cell run: 3487/22355
    20:12 Right HRS magnet settled, reduced beam current to 5 uA and took H2 reference cell run: 3488/22356
    20:55 Multi-foil optics run: 3490/22358
    21:20 Multi-foil optics run: 3491/22359 (raster off)
    21:38 turned BB magnet off, raster off, optics run: 3493/22360
    21:55 BB off raster on optics run 3494/22361, do not have end-of-run file for run 3493. Ended both and restarted run 3495/22362
    22:32 called for escorted access to remove the BB sieve slit
    22:40 BB magnet reset to 518A
    23:00 started Empty reference cell run 3496/22363
    23:20 N2 reference run (19 psig) run 3497/22364
    23:45 Filled in pure 3He gas (135 psig) started run 3498/22365
    New Happex run 31517